OpenStudy (anonymous):

How would you solve this system?: y=2x 5x-y=30

hero (hero):

By substitution

hero (hero):

5x - 2x = 30 3x = 30 x = 30/3 x = 10 y = 20

OpenStudy (anonymous):

Oh so like: 5x-(2x)=30?

hero (hero):

yup